OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − x2 = −11x + 24? (−3, 0) and (8, 0) (3, 0) and (−8, 0) (−3, 0) and (−8, 0) (3, 0) and (8, 0)

hartnn (hartnn):

to get x-intercept, just plug in y=0 in y − x2 = −11x + 24

hartnn (hartnn):

\(y − x^2 = −11x + 24 \\ − x^2 = −11x + 24\) can you find 2 values of x from here ?

OpenStudy (anonymous):

To find the x-intercept(s) of that graph, we have to solve the equation : \[ x^2 −11x + 24=0\\ (x-8)(x-3)=0 \\ x=8\text{ or } x=3 \]